a particle falling under gravity is subject to a retarding force proportional to its velocity.find its position of time , if its starts from rest , and show that it will eventually reach a terminal velocity

1
Expert's answer
2022-04-05T09:45:03-0400

Answer


Force is given by

F=kv

Equilibrium

Mg = kv

"K\frac{ds}{dt}=Mg\\"


By integration

"S=\frac{Mgt}{K}"

"S\alpha t"